1. An electron in a uniform electric field experiences a force of4.0×10^-11.what are the magnitude and direction of E at this point?
1
Expert's answer
2021-10-15T10:22:20-0400

"E = \large\frac{F}{q}" "= \large\frac{4*10^{-11}}{1.6*10^{-19}}" "= 2.5 * 10^8 \large\frac{N}{C}" - the magnitude of E and it is direction the same with electron